Gunnison Company had the following equivalent units schedule and cost information for its Sewing Department for the month of December:

Direct Materials         Conversion Costs
  Units started and completed 49,000 49,000
  Add: Units in beginning work in process ×
         Percentage complete:
         6,000 × 0% direct materials
         6,000 × 50% conversion Costs 3,000
  Add: Units in ending work in process ×
         Percentage complete:
         10,000 × 100% direct materials 10,000
         10,000 × 35% conversion Costs 3,500
  Equivalent units of output 59,000 55,500
  Costs:
         Work in process, December 1:
           Direct Material $78,000
           Conversion Costs 9,000
           Total work in process $87,000
         Current costs:
           Direct Material $826,000
           Conversion Costs 111,000
           Total current costs $937,000

Required:

1. Calculate the unit cost for December, using the FIFO method.
$ per equivalent unit

2. Calculate the cost of goods transferred out, calculate the cost of EWIP, and reconcile the costs assigned with the costs to account for.

Cost of goods transferred out $
Cost of EWIP $
Cost to account for:
BWIP $
Current (December)
  Total $

3. What if you were asked for the unit cost from the month of November? Calculate November's unit cost.
$ per equivalent unit
Answer #1
1 Current cost: Answer
Materials = Direct materials cost/Equivalent unit of output =826000/59000 14.00
Conversion = conversion cost/Equivalent unit of output = 111000/55500 2.00
Cost per equivalent unit TOTAL 16.00
2
Cost of goods transferred out = Total work in process + (Unit started & completed * Cost per Equivalent unit) + (Conversion cost of 50% * conversion per unit) =87000+(49000*16)+(3000*2) 877000
Cost of EWIP =( 100% of direct materials * Cost per Equivalent unit )+ ( 35% conversion cost * Cost per Equivalent unit) =(10000*14)+(3500*2) 147000
Cost to account for:
BWIP 87,000
Current(December) 937000
Total 1,024,000
3
Materials = work in process / percentage balance WIP =78000/5000 15.6
Conversion =9000/2500 3.6
Cost per equivalent unit 19.20